### v2

# Rudyak’s conjecture for lower dimensional 1-connected manifolds

Abstract.
Rudyak’s conjecture states that for any degree one map $f:M\to N$ between oriented closed manifolds there is the inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$ for the Lusternik-Shnirelmann category. We prove the Rudyak’s conjecture for $n$-dimensional simply connected spin manifolds for $n\leq 8$.
2000 Mathematics Subject Classification:
55M30
1. Introduction
The reduced Lusternik-Schnirelmann category (briefly LS-category) $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X$ of a topological space $X$ is the minimal number $n$ such that there is an open cover $\{U_{0},\dots,U_{n}\}$ of $X$ by $n+1$ contractible in $X$ sets. We note that the LS-category is a homotopy invariant. Most of deep applications of the LS-category are based on the classical Lusternik and Schnirelmann theorem [2] which states that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M+1$ gives a low bound for the number of critical points on a manifold $M$ of any smooth not necessarily Morse function. The concept of the LS-category and the above theorem were brought by Lusternik and Schnirelmann to solve the Poincare problem on the existence of three closed geodesics on a 2-sphere for any Riemannian metric on it [14],[15]. In modern time Yu. Rudyak applied the LS-category in his approach to the Arnold conjecture on the number of fixed points of symplectomorphisms [19].
The LS-category is a numerical homotopy invariant which is difficult to compute. Clearly $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(S^{n})=1$. The converse is also true in view of the classical results by Smale, Freedman, and Perelman: If $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ then $M$ is homeomorphic to a sphere. It was proven in [6] that if $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=2$, then the fundamental group of $M$ is free.
Moreover many natural conjectures on the LS-category turned out to be false. Perhaps the most famous was the Ganea Conjecture of the 60s (see [8]) which states that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M\times S^{n})=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M+1$ for a closed manifold $M$. On the break of the 20th century N. Iwase constructed two simply-connected 16-dimensional counter-examples $M_{2}$ and $M_{3}$ to the Ganea Conjecture. His examples are quite delicate and they are based on some phenomena for primary 2-components and primary 3-components in the unstable homotopy groups of spheres.
In the late 90s Rudyak proposed the following [18]:
Conjecture 1.1.
For a degree one map $f:M\to N$ between closed orientable manifolds, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$.
In [18] Rudyak proved his conjecture with some restrictions on dimension, connectivity and the LS-category of the domain. It turned out that his conjecture is quite difficult even in the case of the most simple maps of degree one $f:M\#N\to N$, the collapsing of a summand in the connected sum [7]. Perhaps the most interesting and still a challenging case of Rudyak’s conjecture is the case of birational morphism $f:M^{\prime}\to M$ between nonsingular projective varieties. We note when $M^{\prime}$ is a blowup of $M$ at a point the Rudyak’s conjecture for $f$ follows from the above formula.
Rudyak’s conjecture like many other natural conjectures on the LS-category could be false. In [5] there was an attempt to construct a counterexample to the Rudyak conjecture by means of Iwase’s manifolds. Using Iwase’s examples the first author constructed a map of degree one $f:M\to N$ between 32-dimensional simply-connected manifolds which would be a counter-example to Rudyak’s conjecture if $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{2}\times M_{3})\geq 5$. We recall that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{2})=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{3})=3$. The classical inequality [2]
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(X\times Y)\leq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(X)+\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(Y)$
and computations in [5] imply that $4\leq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{2}\times M_{3})\leq 6$.
A different idea for a counterexample was suggested in [4].
In any case we believe in Rudyak’s conjecture for lower dimensional manifolds. In [20] Rudyak verified his conjecture for manifolds of dimension $\leq 4$. In this paper we prove Rudyak’s conjecture for simply connected spin manifolds of dimension $\leq 8$ (Theorem 4.6). At the end of the paper we give an example of non-spin simply connected 7-dimensional manifold. The existence of such an example does not allow us to extend our approach to non-spin manifolds.
2. Preliminaries
2.1. Properties of maps of degree one
The following is well-known.
Proposition 2.1.
Let $f\colon M\to N$ be a map between closed connected orientable $n$-manifolds of degree one. Then the map $f_{*}\colon H_{*}(M;R)\to H_{*}(N;R)$ is a split epimorphism and $f^{*}\colon H^{*}(N;R)\to H^{*}(M;R)$ is a split monomorphism for any coefficient ring.
Proof.
The spitting map for $f_{*}:H_{*}(M;R)\to H_{*}(N;R)$ is the composition
$H_{*}(N;R)\stackrel{{\scriptstyle PD}}{{\to}}H^{n-*}(N;R)\stackrel{{\scriptstyle f^{*}}}{{\to}}H^{n-*}(M;R)\stackrel{{\scriptstyle PD}}{{\to}}H_{*}(M;R)$
and for $f^{*}:H^{*}(N;R)\to H^{*}(M;R)$ is the composition
$H^{*}(M;R)\stackrel{{\scriptstyle PD}}{{\to}}H_{n-*}(M;R)\stackrel{{\scriptstyle f_{*}}}{{\to}}H_{n-*}(N;R)\stackrel{{\scriptstyle PD}}{{\to}}H^{*}(N;R).$
2.2. G. W. Whitehead Theorem
For nice spaces, such as CW complexes, it is an easy observation that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq\dim X$. D.P. Grossmann [9] and independently G.W. Whitehead [24][2] proved that for simply connected CW complexes $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq\dim X/2$. Moreover, the following inequality holds true [2].
Theorem 2.2.
For $k$-connected CW complexes
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq\dim X/(k+1).$
2.3. Ganea’s approach to the LS-category
Recall that an element of an iterated join $X_{0}*X_{1}*\cdots*X_{n}$ of topological spaces is a formal linear combination $t_{0}x_{0}+\cdots+t_{n}x_{n}$ of points $x_{i}\in X_{i}$ with $\sum t_{i}=1$, $t_{i}\geq 0$, in which all terms of the form $0x_{i}$ are dropped. Given fibrations $f_{i}\colon X_{i}\to Y$ for i=0,,ni=0,...,n, the fiberwise join of spaces $X_{0},...,X_{n}$ is defined to be the space
$X_{0}*_{Y}\cdots*_{Y}X_{n}=\{\ t_{0}x_{0}+\cdots+t_{n}x_{n}\in X_{0}*\cdots*X_{n}\ |\ f_{0}(x_{0})=\cdots=f_{n}(x_{n})\ \}$
and the fiberwise join of fibrations $f_{0},...,f_{n}$ is the fibration
$f_{0}*_{Y}*\cdots*_{Y}f_{n}\colon X_{0}*_{Y}X_{1}*_{Y}\cdots*_{Y}X_{n}\longrightarrow Y$
defined by taking a point $t_{0}x_{0}+\cdots+t_{n}x_{n}$ to $f_{i}(x_{i})$ for any $i$ with $t_{i}>0$. As the name ‘fiberwise join’ suggests, the fiber of the fiberwise join of fibrations is given by the join of fibers of fibrations.
When $X_{i}=X$ and $f_{i}=f:X\to Y$ for all $i$ the fiberwise join of spaces is denoted by $^{n+1}_{Y}X$ and the fiberwise join of fibrations is denoted by $_{Y}^{n+1}f$.
For a topological space $X$, we turn an inclusion of a point $\to X$ into a fibration $p_{0}^{X}:G_{0}(X)\to X$. The $n$-th Ganea space of $X$ is defined to be the space $G_{n}(X)=*_{X}^{n+1}G_{0}(X)$, while the $n$-th Ganea fibration $p_{n}^{X}:G^{n}_{X}\to X$ is the fiberwise join of fibrations $p_{0}^{X}:G_{0}(X)\to X$.
The following theorem was proven by Schwartz in a more general form [23]:
Theorem 2.3.
The inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(X)\leq n$ holds if and only if the fibration $p^{X}_{n}:G_{n}(X)\to X$ admits a section.
We note that the fiber of $p_{0}^{X}$ is homotopy equivalent to the loop space $\Omega(X)$. Thus, the fiber of $p_{n}^{X}$ is homotopy equivalent to the join product $\ast^{n+1}\Omega(X)$.
We recall that the LS-category of a map $f:Y\to X$ is the least integer $k$ such that $Y$ can be covered by $k+1$ open sets $U_{0},\dots,U_{k}$ such that the restrictions $f|_{U_{i}}$ are null-homotopic for all $i$. The above theorem can be extended to maps [4]:
Theorem 2.4.
[…]
We call a map $f:X\to Y$ a $k$-equivalence if it induces isomorphism of homotopy groups in dimension <k<k and an epimorphism in dimension $k$.
The following Proposition was proven in [6] and in [7]:
Proposition 2.5.
Let $f\colon X\to Y$ be an (n2)(n-2)-equivalence of (r1)(r-1)-connected pointed CW-complexes with $r\geq 0$. Then the map $^{s+1}f$ is a (sr+s+n2)(sr+s+n-2)-equivalence.
3. Reduction to the 2 vs 3 case
All manifolds in the paper are considered to be connected, closed, orientable, and smooth.
Proposition 3.1.
Let $N$ be a $n$-dimensional simply-connected manifold, $n\leq 5$. Let $f:M\to N$ be a map of degree 1. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$.
Proof.
By Theorem 2.2 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$. If $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ then $M$ is a homotopy sphere. By Proposition 2.1 NN is a homology sphere. Since $N$ is simply connected, by the Hurewicz theorem it is a homotopy sphere. Then by the J. H. C. Whitehead theorem NN is homotopy equivalent to $S^{n}$. Hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=1$. So, always $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$. ∎
Proposition 3.2.
Let $f:M\to N$ be a map of degree one between simply connected manifolds of dimension $\leq 8$. Suppose that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\geq 3$. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N$.
Proof.
By Theorem 2.2 the LS-category of simply connected manifolds of dimension $\leq 7$ does not exceed 3. So, we have to consider only the case of $n=8$ where $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M,\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 4$. Assume that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N=4$. Then there is no section of the Ganea fibration $p_{3}:G_{3}(N)\to N$. The fiber $F$ of this fibration is homotopy equivalent to a 6-connected space $\ast^{4}\Omega(N)$. The primary obstruction $\kappa\in H^{8}(N;\pi_{7}(F))$ to a section of $p_{3}$ is the only obstruction. Since the primary obstruction is natural, $f^{*}(\kappa)$ is the primary obstruction to the lift of $f$ with respect to $p_{3}$. By Proposition 2.1 $f^{*}(\kappa)\neq 0$. Hence by Proposition 2.4 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f>3$. The inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f\leq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M$ completes the proof. ∎
We recall that if a $n$-manifold $M^{\prime}$ is obtained from $M$ by a surgery then there is a bordism beetween them, i.e. $(n+1)$-dimenisonal manifold $W$ with boundary $\partial W=M\sqcup M^{\prime}$. If the surgery was in dimension $k$, then $W$ is homotopy equivalent to $M$ with some $(k+1)$-cells attached. If $M^{\prime}$ is obtained by a $k$-surgery from $M$, then $M$ is obtained by (nk1)(n-k-1)-surgery from $M^{\prime}$.
Proposition 3.3.
(a) Suppose that $M^{\prime}$ is obtained by a $k$-surgery, $k\leq 4$, from a simply connected $n$-manifold $M$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}W\leq 2$ where $W$ is a bordism of the surgery.
(b) Suppose that $M^{\prime}$ is obtained from a simply connected manifold $M$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ by a surgery in dimension 2. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$.
Proof.
(a) The inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ implies that there is a section $\sigma:M\to G_{2}(M)$ of the Ganea fibration $p_{M}:G_{2}(M)\to M$. The bordism $W$ is homotopy equivalent to the space
$X=M\cup(\cup_{\alpha}D^{k+1}_{\alpha})$
obtained from $M$ by attaching a $(k+1)$-balls along disjoint embedded $k$-spheres $S^{k}_{\alpha}$. We show that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq 2$. Note that the fiber of the Ganea fibration $p_{X}:G_{2}(X)\to X$ is 4-connected. Moreover, for any contractible subset $C\subset X$, the preimage $p_{X}^{-1}(C)$ is 4-connected. Let $p_{X}^{M}:G^{M}_{2}(X)\to M$ be the restriction of $p_{X}$ over $M$. Clearly $p_{X}^{M}$ contains $p_{M}$ as a subfibration. Then the restriction of $\sigma$ to $S_{\alpha}$ is a null-homotopic map $\sigma|_{\alpha}:S^{k}_{\alpha}\to p_{X}^{-1}(D^{k+1})$ for all $\alpha$. Let
$\sigma^{\prime}_{\alpha}:D^{k+1}_{\alpha}\to p_{X}^{-1}(D^{k+1}_{\alpha})$
be an extension. Then the union $\sigma\cup(\cup_{\alpha}\sigma^{\prime}_{\alpha}$ is a homotopy section of $p_{X}$. Then the Homotopy Lifting Property produces an honest section of $p_{X}$.
(b) We note that $W$ is homotopy equivalent to the space
$X^{\prime}=M^{\prime}\cup(\cup_{\alpha}D^{n-2}_{\alpha}).$
Hence the inclusion $j:M^{\prime}\to X^{\prime}$ is (n3)(n-3)-equivalence. Then $\Omega(j):\Omega(M^{\prime})\to\Omega(X^{\prime})$ is an (n4)(n-4)-equivalence. By Proposition 2.5 the induced map
$\ast^{3}\Omega(f):\ast^{3}\Omega(M^{\prime})\to\ast^{3}\Omega X$
between the fibers of $G_{2}(M^{\prime})$ and $G_{2}(X^{\prime})$ is a (2+2+n4)(2+2+n-4)-equivalence. This implies that the map $q:G_{2}(M^{\prime})\to Z$ to the pull-back in
[…]
is a $n$-equivalence. A section $\sigma:X^{\prime}\to G_{2}(X^{\prime})$ defines a section $s:M^{\prime}\to Z$. Since the map $q$ is a $n$-equivalence, the map $s$ can be lifted with respect to $q$. Thus, $p_{M^{\prime}}$ admits a section and, hence, by Theorem 2.3 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$. ∎
4. Main result
The following theorem was proven by I. M. James [13]:
Theorem 4.1.
Let $p\geq 2$ and $X$ be a finite (p1)(p-1)-connected CW complex of dimension $N\leq p(n+1)$. Denote by $\alpha_{X}\in H^{p}(X;\pi_{p}(X))$ the fundamental class of $X$. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(X)=n+1$ if and only if
$0\neq(\alpha_{X})^{n+1}\in H^{p(n+1)}(X;\otimes^{n+1}\pi_{p}(X)),$
where $(\alpha_{X})^{n+1}$ is the $(n+1)$-fold cup product of $\alpha_{X}$.
Proof.
For the proof see [2, Proposition 2.50]. ∎
Theorem 4.2.
Let $N$ be a $3k$-dimensional (k1)(k-1)connected manifold. Let $f:M\to N$ be a map of degree 1. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$.
Proof.
By Theorem 2.2 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)\leq 3$. Let us assume that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=2$. If $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ then $M$ is a homotopy sphere, so $N$ is a homotopy sphere in view of $\deg f=1$, hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=1$, which is a contradiction.
Now let $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N=3$. The case $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ is impossible by previous argument. Suppose that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=2$. Let us assume that $\pi_{k}(N)\neq 0$. Let $u\in H^{k}(N;\pi_{k}(N))$ be the fundamental class. Since $\dim=3k$, we conclude that $u^{3}\neq 0$, by the James Theorem. Since $f^{*}:N\to M$ is a monomorphism (see Proposition 2.1) by previous proposition, we conclude that $(f^{*}(u))^{3}\neq 0$.
Corollary 4.3.
Let $N$ be a $6$-dimensional simply-connected manifold. Let $f:M\to N$ be a map of degree 1. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$.
Lemma 4.4.
Let $f:M\to N$ be a map of degree 1 between $n$-dimensional simply connected manifolds where $n=7$ or $8$ that induces an isomorphism of homotopy groups $f_{*}:\pi_{i}(M)\to\pi_{i}(N)$ for $i\leq n-5$. Then $\max\{\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(f),2\}\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$.
Proof.
[…]
Now, since $M$ and $N$ are simply connected, we have $\pi_{2}(M)$ isomorphic to $H_{2}(M)$ and $\pi_{2}(N)$ is isomorphic to $H_{2}(N)$ by the Hurewicz theorem. Hence, the map $f$ induces an isomorphism $f_{*}:H_{2}(M)\to H_{2}(N)$.
Since $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(f)\leq 2$, By Theorem 2.4 there is a lift $\phi:M\to G_{2}(N)$ of $f$ with respect to the Ganea fibration $p_{2}^{N}$,
[…]
We show that there is a section $s:N\to G_{2}(N)$.
Since $M$ and $N$ are simply connected, the condition of the lemma imply that $f_{*}:H_{i}(M)\to H_{i}(N)=\pi_{i}(N)$ are isomorphisms for $i\leq n-5$ by the classic Whitehead theorem. Then by the Poincare duality, $f^{*}:H^{i}(N)\to H^{i}(M)$ are isomorphisms for $i\geq 5$.
Note that the fiber $F=\ast^{3}\Omega N$ of the fibration $p=p_{2}^{N}:G_{2}(N)\to N$ is is the join product of 3 copies of connected space $\Omega N$ and hence it is $4$-connected. Therefore in the Moore-Postnikov tower for $p:G_{2}(N)\to N$ we have $N=Z_{1}=Z_{2}=Z_{3}=Z_{4}=Z_{5}$:
[…]
We recall that the maps $q_{k}:Z_{k+1}\to Z_{k}$ are principal bundles with fibers $K(\pi_{k}(F),k)$. Thus, there are pull-back diagrams
[…]
where $E_{n}$ are contractible and $\bar{p}_{n}$ are principal bundle with fibers topological abelian groups having homotopy type of the Eilenberg-Maclane spaces $K(\pi_{n-1}(F),n-1)$ (see [16] or [3]).
By commutativity of the above diagrams we obtain
$\kappa_{5}f=\kappa_{5}q_{5}p_{6}\phi=\bar{p}_{5}\kappa_{5}^{\prime}p_{6}\phi.$
Since the map $\bar{p}_{5}\kappa_{5}^{\prime}p_{6}\phi$ factors through a contractible space $E_{6}$, the map $\kappa_{5}f$ is null-homotopic. We recall that $f^{*}$ is injective on the 5-cohomology. Since the map $\kappa_{5}f$ represents the image $f^{*}[\kappa_{5}]$ of the cohomology class represented by $\kappa_{5}:N\to K(\pi_{5}(F),6)$, the map $\kappa_{5}$ is null-homotopic. Therefore, $\kappa_{5}$ admits a lift and, hence, $q_{5}$ admits a section. Since $q_{5}$ is a principal bundle that has a section, it is a trivial bundle over $N$ with the fiber $K(\pi_{5}(F),5)$.
Thus,
$Z_{6}\cong K(\pi_{5}(F),5)\times N.$
The projection $pr:Z_{6}\to K(\pi_{5}(F),5)$ to the fiber defines a cohomology class
$\alpha=[pr\circ p_{6}\circ\phi]\in H^{5}(M;\pi_{5}(F)).$
Since $f^{*}$ is an isomorphism of 5-cohomology, there is $\beta\in H^{5}(N;\pi_{5}(F))$ with $f^{*}(\beta)=\alpha$. The class $\beta$ can be represented by a map
$s:N\to K(\pi_{5}(F),5)$
which defines a section $s_{5}:N\to Z_{6}$ of $q_{5}$ such that the map
$p_{6}\phi=(pr\circ p_{6}\circ\phi,f):M\to K(\pi_{5}(F),5)\times N$
is homotopic to
$s_{5}f=(sf,f):M\to K(\pi_{5}(F),5)\times N.$
Next we show that the map $s_{5}:N\to Z_{6}$ admits a lift with respect to $q_{6}$. This is equivalent to say that
$\kappa_{6}s_{5}:N\to K(\pi_{6}(F),7)$
is null-homotopic. Since $p_{7}\phi$ is a lift of the map $p_{6}\phi$ with respect to $q_{6}$, the composition
$\kappa_{6}p_{6}\phi:M\to K(\pi_{6}(F),7)$
is null-homotopic. Therefore, $\kappa_{6}s_{5}f$ is null-homotopic. This means that $f^{*}$ takes the cohomology class $\kappa_{6}s_{5}]$ to zero. Since by proposition 2.1 the homomorphism $f^{*}$ is injective, we obtain that $\kappa_{6}s_{5}]=0$. Thus, the map $\kappa_{6}s_{5}$ is null-homotopic.
Suppose that $n=7$. Since the map $p_{7}$ is 7-equivalence i.e. it is an isomorphism for $i$-dimensional homotopy groups for $i<7$ and epimorphism for $i=7$, there is a homotopy lift $\sigma:N\to G_{2}(N)$ of the map $s_{6}:N\to Z_{7}$ with respect to $p_{7}$. Then $\sigma$ is a homotopy section of $p$. Since $p$ is a fibration, it admits a regular section. Therefore, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$.
If $n=8$, we construct a lift $s_{7}:N\to Z_{8}$ of $s_{5}$ with respect to $q_{6}q_{7}$ and complete the proof similarly to the case of $n=7$. For that we modify $s_{6}$ to a such a map that $s_{6}f$ becomes homotopic to $p_{7}\phi$.
Since $q_{6}s_{6}=s_{5}$ is an embedding, the map $q_{6}$ is defined by a section of $q_{6}$ on $s_{5}(N)$. Since the map
$\kappa_{6}s_{5}:N\to K(\pi_{7}(F),7)$
is null-homotopic, the restriction of $q_{6}$ over $s_{5}(N)$ is a trivial bundle,
$q_{6}^{-1}(s_{5}(N))\cong s_{5}(N)\times K(\pi_{6}(F),6).$
Since $p_{6}\phi$ is homotopic to $s_{5}f$ and $q_{6}p_{7}\phi=p_{6}\phi$, by the Homotopy Lifting property of $q_{6}$, we can homotop the map $p_{7}\phi$ to a map $\psi:M\to Z_{7}$ such that $q_{6}\psi=s_{5}f$. Thus the map
$\psi:M\to s_{5}(N)\times K(\pi_{6}(F),6)$
has two components $(s_{5}f,\psi_{1})$. Since $f$ induces an isomorphism of 6-dimensional cohomology groups, there is a map
$\xi:N\to K(\pi_{6}(F),6)$
such that $f^{*}$ takes the cohomology class $\xi]\in H^{6}(N;\pi_{6}(F))$ to the cohomology class $\psi_{1}]\in H^{6}(M;\pi_{6}(F))$. Thus, $\psi_{1}$ is homotopic to $\xi f$. We consider a section
$s^{\prime}:s_{5}(N)\cong N\to q_{6}^{-1}(s_{5}(N))\cong s_{5}(N)\times K(\pi_{6}(F),6)$
defined by $\xi$. Then $\psi$ is homotopic to the map $s^{\prime}s_{5}f$. Since $p_{8}\phi$ is a lift of the map $p_{7}\phi$ with respect to $q_{7}$, the composition
$\kappa_{7}p_{7}\phi:M\to K(\pi_{6}(F),7)$
is null-homotopic. Hence $\kappa_{7}\psi$ is null-homotopic and the map $\kappa_{7}s^{\prime}s_{5}f$ is null-homotopic. This means that $f^{*}$ takes the cohomology class $\kappa_{7}s^{\prime}s_{5}]$ to zero. Since the homomorphism $f^{*}$ is injective, we obtain that $\kappa_{7}s^{\prime}s_{5}]=0$. Thus, the map $\kappa_{7}s^{\prime}s_{5}$ is null-homotopic and, hence, there is a lift $s_{7}:N\to Z_{8}$ of $s^{\prime}s_{5}$ with respect to $q_{7}$.
Since the map $p_{8}$ is 8-equivalence, there is a homotopy lift $\sigma:N\to G_{2}(N)$ of the map $s_{7}:N\to Z_{8}$ with respect to $p_{8}$. Then $\sigma$ is a homotopy section of $p$. Since $p$ is a fibration, it admits a regular section. Therefore, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$. ∎
Proposition 4.5.
A degree one map $f:M\to N$ of a $n$-dimensional, $n=7,8$ simply connected spin manifold with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ is bordant by means of surgery in dimension 2 and 3 to a a degree one map $f^{\prime}:M^{\prime}\to N$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f^{\prime}\leq 2$ that induces isomorphisms $f^{\prime}:\pi_{i}(M^{\prime})\to\pi_{i}(N)$ for $i\leq n-5$.
Proof.
For $n=7,8$ by Proposition 2.1 the induced homomorphism $f_{*}:H_{2}(M)\to H_{2}(N)$ is surjective with a finitely generated kernel $K=kerf_{*}$ which is by the Hurewicz Theorem isomorphic to $ker\{f_{*}:\pi_{2}(M)\to\pi_{2}(N)\}$. Let $S_{1},\dots,S_{k}$ be disjoint 2-spheres smoothly embedded in $M$ that generate $K$. For each sphere $S_{i}$ its normal bundle $\nu_{i}$ is classified by a map $\nu_{i}:S_{i}\to BSO(n-2)$. The map $\nu_{i}$ is null-homotopic, since otherwise the evaluation of the Stiefel-Whitney class $w_{2}\in H^{2}(BSO(n-2);\mathbb{Z}_{2})$ on the image $(\nu_{i})_{*}([S_{i}])\in H_{2}(BSO(n-2);\mathbb{Z}_{2})$ would be nonzero. The latter would contradict to the spin condition on $M$. Thus, we can perform a surgery to obtain a bordism $W$ of $M$ to $M^{\prime}$ together with a map $\bar{f}:W\to N$. The restriction $f^{\prime}:M^{\prime}\to N$ of $\bar{f}$ to $M^{\prime}$ has degree one and $f^{\prime}$ induces an isomorphism of 2-dimensional homotopy groups. For $n=7$ by Proposition 3.3 (b) $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$ and, hence, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f^{\prime}\leq 2$.
When $n=8$ by the above we may assume that $f:M\to N$ induces an isomorphism of 2-dimensional homotopy groups. Since $\pi_{3}(BSO(n-3))=0$, we obtain that the normal bundle of any embedded 3-sphere in $M^{\prime}$ is trivial. Hence we can perform surgery on generators of the kernel $ker\{f_{*}:\pi_{3}(M^{\prime})\to\pi_{3}(N)\}$ to get a bordism $W$ between $M$ and $M^{\prime}$ together with a map $\hat{f}:W\to N$ that agrees with $f$ on $M$ such that the restriction $f^{\prime}:M^{\prime}\to N$ of $\hat{f}$ to $M^{\prime}$ has degree one and it induces isomorphisms of homotopy groups of $\dim\leq 3$. By Proposition 3.3 (a) we obtain that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}W\leq 2$. Hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}\hat{f}\leq 2$. By Theorem 2.4 $\hat{f}$ admits a lift $\hat{\phi}:W\to G_{2}(N)$ of $\hat{f}$ with respect to $p_{2}^{N}$. The restriction of $\hat{\phi}$ to $M^{\prime}$ is a lift of $f^{\prime}:M^{\prime}\to N$ with respect to $p^{N}_{2}$. Then by Theorem 2.4 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f^{\prime}\leq 2$. ∎
Theorem 4.6.
A degree one map $f:M\to N$ between simply connected spin $n$-manifold for $n\leq 8$ satisfies the Rudyak conjecture,
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N.$
The spin condition can be dropped when $n\leq 6$.
Proof.
When $n\leq 5$ the statement is proven in Proposition 3.1. In view of Proposition 3.2 it suffice to show that there is no degree one map $f:M\to N$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ and $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\geq 3$. The case of $n=6$ is taken care of in Corollary 4.3.
We consider the case when $n\geq 7$ and $M$ is spin. Let $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M=2$. In view of Proposition 4.5 and Proposition 3.3 we may assume that $f$ induces isomorphisms of homotopy groups of dimension $\leq 3$ and $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f\leq 2$. Then by Lemma 4.4 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$. ∎
Remark 4.7.
There are non-spin 7-manifolds with the LS-category 2. We recall that the Wu manifold SU(3)/SO(3) is simply connected and non-spin. A 7-manifold $M$ constructed from the product of a 2-sphere and the Wu manifold
$S^{2}\times(SU(3)/SO(3))$
by a 2-surgery on the factor $S^{2}$ is called a gyration of the Wu manifold [25]. It is simply connected and non-spin. Below we compute its LS-category.
Proposition 4.8.
The LS-category of the gyration of the Wu manifold is 2.
Proof.
Let $Y=SU(3)/SO(3)$. By Theorem 2.2 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}Y=2$. The bordism $X$ defined by the 2-surgery on $Y\times S^{2}$ is homotopy equivalent to the union $(Y\times S^{2})\cup(\{y_{0}\}\times D^{3})$ which is homotopy equivalent to the half-smash product
$Y\rtimes S^{2}=(Y\times S^{2})/(\{y_{0}\}\times D^{3}).$
It is known that for the half-smash product [22]
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(A\rtimes B)=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}A.$
Hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}Y=2$. Let $M$ be the gyration of the Wu manifold. Then $X$ is homotopy equivalent to $X^{\prime}=M\cup D^{5}$. The rest of the proof coincides with the end of the proof of Proposition 3.3 (b). Since the inclusion $j:M\to X^{\prime}$ is a 4-equivalence, $\Omega(j):\Omega(M)\to\Omega(X^{\prime})$ is a 3-equivalence. By Proposition 2.5
$\ast^{3}\Omega(j):\ast^{3}\Omega(M)\to\ast^{3}\Omega(X^{\prime})$
is a $(2+2+3=7)$-equivalence. This implies that the map $q:G_{2}(M)\to Z$ to the pull-back in the diagram
[…]
is a $7$-equivalence. Since $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X^{\prime}=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X=2$, there is a section $\sigma:X^{\prime}\to G_{2}(X^{\prime})$ which defines a section $s:M\to Z$ of $p^{\prime}$. Since the map $q$ is a $7$-equivalence, $s$ can be lifted with respect to $q$. Thus, $p_{M}$ admits a section and, hence, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$.
Since $M$ is not homeomorphic to a sphere, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M>1$ and we obtain that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M=2$. ∎